My Verbal Equation Is To Find Out What:
Twelve less than 2 times a number is equal to 15 minus 7 times the number is.
Any help with this problem would be greatly appreciated:D
,Kia Jayy

Answer by checkley75(3666)

2X-12=15-7X
2X+7X=15+12
9X=27
X=27/9
X=3 ANSWER.
PROOF
2*3-12=15-7*3
6-12=15-21
-6=-6